# Word problems and embedding-obstructions in cellular automata groups on groups

### v2

@@ 2.2 Groups
2.3 Word problems
Now we describe our word problem conventions. We will work with word problems of groups that are not finitely-generated, but are mainly interested in their finitely-generated subgroups. Thus, we use a non-standard convention for word problems. We define $\mathrm{WPs}(G)$ to be the set of word problems of its finitely-generated subgroups. This is a family of languages, and we next define our conventions for complexity theoretic notions of language families.
If $C$ is a complexity class of languages defined directly by resource bounds (such as the classes P, co-NP, PSPACE or co-NEXPSPACE), then for a family of languages $\mathcal{L}$, we say that $\mathcal{L}$ is in $C$ if every language in $\mathcal{L}$ is in $C$. We say $\mathcal{L}$ is $C$-hard if some language in $\mathcal{L}$ is $C$-hard (for polynomial time many-one reductions). We say $\mathcal{L}$ is $C$-complete if $\mathcal{L}$ is in $C$ and is $C$-hard.
Note that $D$-hard (and $D$-complete) can themselves be called complexity classes, and we do not give a definition of “resource bounded” that excludes such classes. Thus saying a language class $\mathcal{L}$ is $D$-hard can seem ambiguous. This will not cause problems; simply do not think of $D$-hard and $D$-complete as complexity classes, but instead derived notions that are defined differently for languages and language families.
@@ 2.5 $\mathrm{PAut}$
  • symbol permutations $\eta_{\pi}$ for $\pi\in\mathrm{Sym}(A)$, which on input $x\in A^{G}$ are defined by $f_{\pi}(x)_{h}=\pi(x_{h})$ for $h\in G$.
@@
This is entirely analogous to [31, Lemma 2.1]. The point is the change between the groups for different choices is in the partial shifts. Since the prime decomposition is unique, these are conjugated by a symbol permutation, which is in the group in any case. ∎
3 Groups with free submonoids
In this section, we prove PSPACE-hardness of the word problem for full shifts with suitable alphabets on groups that contain free submonoids on at least two generators.
Theorem 3.1.
@@ 3 Groups with free submonoids
If $G$ contains a free monoid on two generators, it contains a free monoid on three generators, say powers of $a,b,c\in G$ satisfy no nontrivial positive relations.
We will use alphabet $A=\mathbb{Z}_{7}\times B\times\{0,1,2\}$ for our f.g. group, where $B=Q\times S$. Here, $Q$ with $|Q|=t$, and $S$ with $|S|=s$ are respectively the common sets of states and tape symbols of the nondeterministic Turing machines $T_{1},T_{2}$ from Lemma A.3.
Our reduction takes a word $u_{\mathrm{init}}$ of length $n=2^{k}$ (note that one should think of this as a polynomial size input whose size just happens to be of nice form, rather than an exponential length word), and constructs an automorphism that performs a permutation $\pi$ on the $\mathbb{Z}_{7}$-component of $A$ whenever it sees a tree of the form described in Lemma A.3 on the configuration, encoded in words read on the $B$-components, and with $q_{\mathrm{init}}u_{\mathrm{init}}\in QS^{n}$ as the initial configuration on the tree (where $q_{\mathrm{init}}$ is the common initial state of $T_{1},T_{2}$).
As in the lemma, the computation steps in the tree are given by the Turing machines $T_{0},T_{1}$ with common initial state $q_{\mathrm{init}}$, states $Q$ and tape alphabet $S$. Note that we now have the state serially next to tape symbols rather than use a Cartesian product, just for notational simplicity. The tree still has depth $n$, but the Turing machine configurations on it have length $n+1$.
This will indeed many-one reduce the language $L$ from Lemma A.3 to the word problem of the automorphism group of $A^{G}$, since the automorphism corresponding to the word $u_{\mathrm{init}}$ will be nontrivial if and only if the permutation is ever performed, and it is performed if and only if a tree proving $u_{\mathrm{init}}\in L$ exists.
The main difficulty here is that the tree has exponential size in $|u_{\mathrm{init}}|$, so we cannot use Barrington’s theorem to directly verify the shape of the tree (which is essentially all we had to do in the $\mathbb{Z}$-case in [29]). The computation will be “initialized” with Barrington’s theorem, and in addition we will use a tree propagation argument to efficiently send a permutation down the computation tree in order to check that the computation is correct everywhere.
We want to use the following generators, which are obviously in $\mathrm{Aut}(A^{G})$:
  • partial right shifts of the first track, i.e. writing $x\in A^{G}$ as $(x_{1},x_{2})\in\mathbb{Z}_{7}^{G}\times(\mathbb{Z}_{m}\times\{0,1\})^{G}$, we take the map $\hat{g}(x_{1},x_{2})_{c}=((x_{1})_{cg},(x_{2})_{c})$ for all $g\in G$.
  • (even) symbol permutations affecting only $\mathbb{Z}_{7}$: for each $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ and each $s\in\mathbb{Z}_{m}\times\{0,1\}$, and writing $x\in A^{G}$ as $(x_{1},x_{2})$ as above, we take the permutation $\alpha_{\pi|s}$, defined by
    $\alpha_{\pi|s}(x_{1},x_{2})_{c}=\begin{cases}(\pi(x_{1})_{c},(x_{2})_{c})&\mbox{ if }(x_{2})_{c}=s\\ ((x_{1})_{c},(x_{2})_{c})&\mbox{ otherwise.}\end{cases}$
  • for each permutation $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$, $s\in\mathbb{Z}_{7}$ not in the support of $\pi$ and $g\in G\setminus\{1\}$, the map $\phi=\phi_{\pi,s,g}$ defined by
    $\phi(x_{1},x_{2})_{c}=\begin{cases}(\pi((x_{1})_{c}),(x_{2})_{c})&\mbox{ if }(x_{1})_{cg}=s\\ ((x_{1})_{c},(x_{2})_{c})&\mbox{ otherwise.}\end{cases}$
Note that the generators only affect the first track (though the way it is changed can depend on the latter tracks). Call this group $\mathcal{G}$.
We show that $\mathcal{G}$ is a subgroup of $\mathrm{PAut}=\mathrm{PAut}(\mathbb{Z}_{7},B,\{0,1\})$, as required in the statement. This is only nontrivial for $\phi_{\pi,s,g}$. To show it is in $\mathrm{PAut}$, we will use the $\{0,1,2\}$-component of the state to “move” a permutation from one cell to another. The $\{0,1,2\}$-component acts here as a borrowed bit in the terminology [37] (or perhaps borrowed trit would be more accurate); this means that while it is used in the construction, its initial value is irrelevant and this value is returned at the end.
We now give the detailed proof of $\phi_{\pi,s,g}\in\mathrm{PAut}$. Write $\pi=[\pi_{0},\pi_{1}]$ where neither $\pi_{i}$ has $s$ in its support. We define four automorphisms by describing their effect on the symbol $x_{c}$ of the input $x\in A^{G}$, for $c\in G$. Define an automorphism $\theta_{b}$ that, if the $\{0,1,2\}$-component of the alphabet at $x_{c}$ is $b\in\{0,1\}$, permutes the first component of $x_{c}$ by $\pi_{i}$. Clearly this is in $\mathrm{PAut}$. Then define an automorphism $\theta$ that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) if the $\mathbb{Z}_{7}$-component of $x_{cg}$ is $s$. This is conjugate to a symbol permutation by a partial shift, so again in $\mathrm{PAut}$. Finally, let $\theta^{\prime}$ be the automorphism that permutes the $\mathbb{Z}_{3}$-component of $x_{c}$ by the rotation (0 1 2)012(0\;1\;2)( 0 1 2 ) in any case. Note that
@@
From now on, let us write configurations $x\in A^{G}$ as $x=(x_{1},x_{2},x_{3})\in\mathbb{Z}_{7}^{G}\times\mathbb{Z}_{m}^{G}\times\{0,1\}^{G}$. The TM configuration at $g$ is $i\mapsto x_{2}|gc^{i}$ for $i\in\{0,\ldots,n\}$. Note that is a word in $(S\times Q)^{n+1}$, and might not actually be a TM configuration in the usual sense (as it might contain multiple states, or none). We say a cell $g$ is TM-good if the TM configuration at $g$ has exactly one symbol in $Q$.
The index at $g$ is $i\mapsto x_{3}|gc^{i}$ for $i\in\{0,\ldots,n-1\}$. It is a ternary word of length $n$ over alphabet $\{0,1,2\}$, which we interpret as a index to the tree witness from Lemma A.3. The symbol $2$ is used for padding the words. If the index at $g$ is $uv$ where $u\in\{0,1\}^{*}$ and $v\in 2^{*}$, then the counter value at $g$ is $|u|$. Note that the counter value may not exist for all $g$.
(As an aside, in the present proof, only the counter value at $g$ really matters in the following, so we could equally well use 00 in place of 00 and $1$, and $1$ in place of $2$, but above it was useful to have three states in this component, and in later variants of this argument we will have more uses for the precise indices.)
We say $g$ is good (with respect to a configuration $x\in A^{G}$) if
  • $g,ga,gb$ are TM-good,
  • the TM-configurations $g$ and $ga$ are in the transition relation of $T_{1}$, and the ones at $g$ and $gb$ are in the transition relation of $T_{2}$,
  • if the index at $g$ is in $u2^{*}$ for $u\in\{0,1\}^{*}$, the indices at $ga,gb$ are in $u02^{*},u12^{*}$ respectively,
  • if the counter value at $g$ is 00, then the TM configuration is precisely $q_{0}u_{\mathrm{init}}$.
@@
Let $\gamma_{\pi,\ell}$ be the automorphism that permutes $\mathbb{Z}_{7}$ at $g$ by $\pi$ if and only if the counter value is exactly $\ell$. Again, one easily checks that there is a logarithmic depth formula for comparing a number to a constant, and uses Theorem A.2.
Next, we perform what we call dyadically accelerated funneling which we shorten as DAF. In such a proof, a permutation will intuitively travel down a “funnel” of depth $n$. In the present proof, the funnel is a full binary tree, but we also use other kinds of trees (in Theorem 5.1) and discrete analogs of Euclidean cones (Theorem 4.1) in other proofs. [^1] We introduce a term (DAF) for this (informal!) concept, so we can refer back to this idea in other proofs, and since we consider this the main new idea of the paper as far as constructions go. To explain the terminology a little, our mental picture of “funneling” is that “verification checks” travel down a funnel in the form of controlled permutations in intermediate automorphisms in the construction, and finally a permutation is extracted at the bottom. Then the point is that by an induction on dyadic intervals, we can perform such a funneling process in logarithmically many steps.
For $i\in[0,k]$ and $j\in[0,2^{k-i}-1]$, consider now an interval $I=[j2^{i},(j+1)2^{i}]\subset[0,n]$, let $s\in\mathbb{Z}_{7}$ and $\pi\in\mathrm{Alt}(\mathbb{Z}_{7})$ is a $3$-rotation. Note that $|I|\geq 2$ always. We define an automorphism $\psi_{j,i,s,\pi}$ as follows: The latter two tracks never change their values. On the first track, in position $c\in G$, the value changes by an application of $\pi$ if and only if the following hold:
  • the counter value at $c$ (exists and) is equal to $j2^{i}$;
  • for all $\ell\in[0,2^{i}-1]$ and any $h\in\{a,b\}^{\ell}$, the node $ch$ is good;
  • for all $\ell\in[1,2^{i}-1]$ and any $h\in\{a,b\}^{\ell}$, the $\mathbb{Z}_{7}$-value at $ch$ is 00;
  • for all $h\in\{a,b\}^{2^{i}}$, the $\mathbb{Z}_{7}$-value at $ch$ is $s$.
Note that the counter value at $ch$ is $j2^{i}+\ell$ for all $\ell\in[0,2^{i}]$ and any $h\in\{a,b\}^{\ell}$ by the first two items.
The DAF argument is specifically the proof (method) showing that the automorphisms $\psi_{0,k,s,\pi}$ have polynomial word norm in $n$ in the group $\mathcal{G}$. Note that this is not immediate from Barrington’s theorem, from which for checking conditions on a tree with size exponential in $n$, we can only conclude a word norm polynomial in $2^{n}$.
Suppose first that $s$ is not in the support of $\pi$. Then for $i=0$ and $j\in[0,2^{k-i}-1]$ we can build the above automorphisms from the automorphisms $\beta_{\pi}$, $\phi_{\pi,s,g}$ and $\gamma_{\pi,\ell}$ described above. Specifically, let $\pi=[\pi_{0},\pi_{1},\pi_{2},\pi_{3}]$ where the $\pi_{i}$ are $3$-rotations not having $s$ in their support. We claim that
$\psi_{j,0,s,\pi}=[\beta_{\pi_{0}},\phi_{\pi_{1},s,a},\phi_{\pi_{2},s,b},\gamma_{\pi_{3},j}].$
The norm is clearly polynomial in $n$.
As for correctness of the formula, the logic is similar to Lemma A.9. Namely, the four automorphisms that appear in the commutator only change the $\mathbb{Z}_{7}$-values, and do not change the positions where the $s$ appears in the $\mathbb{Z}_{7}$-component never change. Thus in each cell, the permutation $\pi_{i}$ (or its inverse) either applies each time the corresponding automorphism is applied, or never applies.
We conclude that the only way their effect does not cancel in a particular node is that all of them perform the permutation in that node $h$. That $\beta_{\pi_{0}}$ is performed means $h$ is good. That $\phi_{\pi_{1},s,a}$ and $\phi_{\pi_{1},s,b}$ perform the permutation means $ha$ and $hb$ both contain $s$ in the $\mathbb{Z}_{7}$-component, and that $\gamma_{\pi_{3},j}$ is performed means that the counter value is $j$ at $h$. This is precisely the set of conditions under which $\pi=[\pi_{0},\pi_{1},\pi_{2},\pi_{3}]$ is supposed to be performed, proving the claimed equality for $i=0$ and $s$ not in the support of $\pi$.
@@
Let $N_{0}$ be an upper bound on the norms of the automorphisms $\psi_{j,0,s,\pi}$, which we have now concluded to be polynomial in $n$, say $O(n^{t})$.
The crux of DAF is to perform induction to get $\psi_{j,i,s,\pi}$ for larger $i$, with only exponential blow-up as $i$ is increased. We have
$[j2^{i},(j+1)2^{i}]=[2j2^{i-1},(2j+1)2^{i-1}]\cup[(2j+1)2^{i-1},2(j+1)2^{i-1}]$
and the two intervals have intersection $\{(2j+1)2^{i-1}\}$. We claim that
@@
where $\pi=[\pi_{1},\pi_{2}]$, and $\pi_{3}$ is any permutation taking 00 to $1$.
@@
Suppose then that $\psi_{j,i,s,\pi}$ does not apply the permutation. There are several reasons why this might not happen:
  • we may have that for some $\ell\in[0,2^{i}]$ and some $h\in\{a,b\}^{\ell}$, the node $ch$ does not have counter value $j2^{i}+\ell$; or
  • for some $\ell\in[1,2^{i}-1]$ and some $h\in\{a,b\}^{\ell}$, the $\mathbb{Z}_{7}$-value at $ch$ is not 00; or
  • for some $h\in\{a,b\}^{2^{i}}$, the $\mathbb{Z}_{7}$-value at $ch$ is not $s$.
We look at these three cases one by one. In the first case, for some $\ell\in[0,2^{i}]$ and some $h\in\{a,b\}^{\ell}$, the node $ch$ does not have counter value $j2^{i}+\ell$. Since none of the three automorphisms in the right-hand-side formula change the counter values, this will stay true at $c$ at all the applications. Thus, depending on whether $\ell\in[0,2^{i-1}]$ or $\ell\in[2^{i-1},2^{i}]$, either the maps $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ will not apply, or the conjugating map $\psi_{2j+1,i-1,s,\pi_{3}}$ will not apply. In the former case, nothing happens, and in the latter case, $\psi_{2j+1,i-1,s,\pi_{3}}$ does not flip one of the values in $c\{a,b\}^{2^{i-1}}$ meaning one of $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ ends up not applying a permutation at $c$ (because for them both to apply, the entire frontier must flip its value), which leads to the permutation cancelling at $c$.
Suppose then that we are not in the first case, so the counter value at $ch$ is equal to $j2^{i}+\ell$ whenever $\ell\in[0,2^{i}]$ and $h\in\{a,b\}^{\ell}$. Suppose that we are, however, in the third case, so for some $h\in\{a,b\}^{2^{i}}$, the $\mathbb{Z}_{7}$-value at $ch$ is not $s$. Then, $\psi_{2j+1,i-1,s,\pi_{3}}$ never applies, and we are in a similar case as above.
Suppose then that we are not in the first or third case, but that we are in the second case. So for some $\ell\in[1,2^{i}-1]$ and some $h\in\{a,b\}^{\ell}$, the $\mathbb{Z}_{7}$-value at $ch$ is not 00. If this happens for some $\ell\neq 2^{i-1}$, we are in a similar situation as in the previous case: either $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ will not apply, or the conjugating map $\psi_{2j+1,i-1,s,\pi_{3}}$ will not apply, and everything cancels at $c$.
If this only happens with $\ell=2^{i-1}$, everything is as in the case where $\psi_{j,i,s,\pi}$ would normally apply, except that at exactly depth $2^{i-1}$, some values are not necessarily 00. The conjugating map $\psi_{2j+1,i-1,s,\pi_{3}}$ now always applies. Nevertheless, again one of $\psi_{2j,i-1,0,\pi_{1}},\psi_{2j,i-1,1,\pi_{2}}$ ends up not applying a permutation at $c$.
This case analysis concludes the proof of Equation 2.
The word norm of such an element at level $i$ is now $N_{i}\leq 8N_{i-1}$, so $N_{k}\leq 8^{\log_{2}2^{k}}=8^{k+1}N_{0}=O(n^{3}n^{t})$.
Thus, $\psi_{0,k,s,\pi}$ can be presented as a word over the generators of $\mathrm{PAut}$ with word norm polynomial in $n$ (which in turn, as we recall, is the length of the initial configuration $u_{\mathrm{init}}$). As discussed in the beginning of the proof, this element is nontrivial if and only if there is a way to fill a tree with a computation of $T_{1}$ and $T_{2}$ starting from $q_{0}u_{\mathrm{init}}$, concluding the reduction from $L$ to the word problem of $G$ (thus $\mathrm{PAut}$). ∎
4 Unconditional nonembedding results

### v3

@@ 2.2 Groups
The word problem of a group $G$, with respect to a finite generating set $S$, is the language $\{w\in S^{*}\;|\;w=_{G}1_{G}\}$, where $=_{G}$ denotes equality when evaluated in the group $G$.
@@ 2.4 $\mathrm{PAut}$
  • symbol permutations $\eta_{\pi}$ for $\pi\in\mathrm{Sym}(A)$, which on input $x\in A^{G}$ are defined by $\eta_{\pi}(x)_{h}=\pi(x_{h})$ for $h\in G$.
@@
This is entirely analogous to [30, Lemma 2.1]. The point is the change between the groups for different choices is in the partial shifts. Since the prime decomposition is unique, these are conjugated by a symbol permutation, which is in the group in any case. ∎
Suppose $A=C\times B$, $G$ is a finitely-generated group, $\pi\in\mathrm{Sym}(C)$, $c\in C$ is not in the support of $\pi$, and $s\in G\setminus\{1_{G}\}$. Then we define $\phi_{\pi,c,s}$ as the automorphism of $A^{G}$ that applies $\pi$ to the $C$-component of the symbol at $g$ if the $C$-symbol at $gs$ is equal to $c$.
Lemma 2.5.
Let $|C|\geq 4,|B|\geq 2$ be alphabets and $\pi\in\mathrm{Alt}(C)$. Suppose $c$ is not in the support of $\pi$. Then $\phi_{\pi,c,s}\in\mathrm{PAut}_{G}(B\times C)$.
@@
Assume $B=\mathbb{Z}_{h}$ for some $h\geq 2$. Use Ore’s theorem to obtain for each $\pi\in\mathrm{Alt}(C)$ a decomposition $\pi=[\pi_{0},\pi_{1}]$ where neither $\pi_{i}$ has $c$ in its support. We define automorphisms by describing their effect on the symbol $x_{g}$ of the input $x\in A^{G}$, for $g\in G$. Define an automorphism $\theta_{b}$ that, if the $B$-value at $x_{gs}$ is $b$, permutes the first component of $x_{g}$ by $\pi_{b}$. Clearly this is in $\mathrm{PAut}$. Next define an automorphism $\theta$ that permutes the $B$-component of $x_{g}$ by (0 1)01(0\;1)( 0 1 ) if the $C$-component of $x_{g}$ is $c$. This is conjugate to a symbol permutation by a partial shift, so again in $\mathrm{PAut}$. Finally, let $\theta^{\prime}$ be the automorphism that permutes the $B$-component of $x_{g}$ by a full length $|B|$ rotation in any case.
@@
For finite $S\Subset G$, define $\phi_{\pi,c,S}$ as the automorphism of $A^{G}$ that applies $\pi$ to the $C$-component of the symbol at $g$ if the $C$-symbol at $gs$ is equal to $c$ for all $s\in S$.
Lemma 2.6.
Let $S\Subset G$ be finite, let $|C|\geq 6,|B|\geq 2$ be alphabets and $\pi\in\mathrm{Alt}(C)$. Suppose $c$ is not in the support of $\pi$. Then $\phi_{\pi,c,S}\in\mathrm{PAut}_{G}(B\times C)$.
Proof.
We may write $S=\{s_{1},s_{2},\ldots,s_{\ell}\}$. Since $|C|\geq 6$, we may write $\pi=[\pi_{1},\pi_{2},\ldots,\pi_{\ell}]$ where none of the $\pi_{i}\in\mathrm{Alt}(C)$ have $c$ in their support. Then
$\phi_{\pi,c,S}=[\phi_{\pi_{1},c,S},\phi_{\pi_{2},c,S},\ldots,\phi_{\pi_{\ell},c,S}]$
since the condition for some $\phi_{\pi_{i},c,S}$ (equivalently, all of them) applying the permutation is only a function of which nodes the symbol $c$ appears as the $C$-value, which is not affected by the permutations $\pi_{i}$. ∎
3 Ripple catching
In this section, we present the main new technology of the paper, namely ripple catching. This is a simple trick that allows us to use parallelism to “check” that a large cone exists at a position, in the positive direction in a graph, and that certain successor relations (which can be anything checkable in NC1) are respected. Specifically, a particular permutation is applied at a node if and only if the cone rooted at that element exists and everywhere has locally valid content, in some desired sense.
To realize this, we consider a “ripple of permutations” down from the rim of the cone to its root, and we condition a permutation at the root of the cone on the ripple reaching the it. Using a commutator of two such ripples, we can ensure that only a “complete” ripple results in a nontrivial permutation.
Specifically, we will consider a graph whose nodes have colors in $C\times B$, and we will perform a permutation on the $C$-component, controlled by an essentially arbitrary condition on the colors $B$ (defined by some successor relation), and also the condition that $C$-colors are all 00 outside the root of the cone.
Definition 3.1.
@@ 3 Ripple catching
In the present section, we will not need $s$-edges for $s\in S^{-1}$, but we need them in Section 6 so we include them in this definition.
@@
Let $C$ and $S$ be finite sets, let $B$ be a countable set, and suppose we have associated to each $s\in S$ a successor relation $R_{s}\subset B\times B$. If $(b,b^{\prime})\in R_{s}$, we say $b^{\prime}$ is an $s$-successor of $b$. Often, the successor relation is deterministic, in that there is at most one $s$-successor for each color $b\in B$. In this case, we can also describe the successor relations by giving a partial action of $S$ (and thus $S^{*}$).
Further suppose $B$ is ranked meaning there is a rank function $\rho:B\to\mathbb{N}\cup\{\bot\}$ such that $\rho(b^{\prime})=\rho(b)+1$ or $\rho(b^{\prime})=\bot$ for all $b,b^{\prime}\in B$ such that $b^{\prime}$ is an $s$-successor of $b$ for some $s\in S$. Consider the set $\mathcal{Q}_{S,C,B}$ of all $S$-labeled $(C\times B)$-colored good graphs on a fixed vertex set. [^1] The fixed vertex set is just for set-theoretic reasons, we may assume it is large enough for all practical purposes. Since the group is countable, there is in any case a faithful action on a countable subset.
Say a node $u$ in a graph is successful if its color is $(c,b)$, and for each $s\in S$, its $s$-successor $v$ exists, and the $B$-component of the color of $v$ is an $s$-successor of $b$ for each $s$.
Definition 3.2.
Let $S,C,B$ be finite sets, with $|C|\geq 6$. Define the ripple group $\mathcal{G}_{S,C,B}$ by the following generators acting on $\mathcal{Q}_{S,C,B}$:
  • If $\pi\in\mathrm{Alt}(C)$, $s\in S$, and $c\in C$ is not in the support of $\pi$, the generator $\phi_{\pi,c}$ applies $\pi$ on the $C$-component $u$ if whenever $us$ exists, the $C$-color of $us$ is $c$.
  • If $\pi\in\mathrm{Alt}(C)$, the generator $\gamma_{\pi}$ applies $\pi$ on the $C$-component of the color of a node $u$, if $u$ is successful.
  • If $\pi\in\mathrm{Alt}(C)$ and $n\in\mathbb{N}$, the generator $\beta_{\pi,\ell}$ applies $\pi$ on the $C$-component of the color of a node $u$ if the $B$-component of $u$ contains $b$ such that $\rho(b)=\ell$.
For each $n$, the group $\mathcal{G}_{S,C,B}$ is endowed with the left-invariant word $n$-metric given by these generators, with $\ell$ bounded from above by $n$.
Some simple remarks are in order:
  1. 1.
    None of the generators affect the shape of the graph or the $B$-components of colors, only the $C$-components.
  2. 2.
    All of $\phi_{\pi,c},\gamma_{\pi},\beta_{\pi,\ell}$ obviously define bijections: they perform permutations on $C$-components of some nodes, and the conditions under which these permutations are not affected by the automorphism itself, so they have finite order.
  3. 3.
    The name of $\phi_{\pi,c}$ is very close to $\phi_{\pi,c,s}$ and $\phi_{\pi,c,S}$ from the previous section, but there is no direct clash, and $\phi_{\pi,c}$ can be seen as a generalization (or abstraction) of $\phi_{\pi,c,S}$.
  4. 4.
    The word $n$-metric is proper for all $n$ in the sense that balls are finite, but since these are not a generating set for the entire group $\mathcal{G}_{S,C,B}$, distances can be infinite (though one can alternatively consider the metric in the group generated by the finitely many $n$-bounded generators).
We now define an important element of this group, which “combines” the three types of generators into one.
Definition 3.3.
Let $S,C,B$ be finite sets, with $|C|\geq 6$. If $\pi\in\mathrm{Alt}(C)$, then the map $\psi_{\pi,\ell,c}$ acts on $\mathcal{Q}_{S,C,B}$ as follows: it applies $\pi$ on the $C$-component of a node $u$ if
  • $u$ is successful,
  • the $C$-component of the color of $us$ is $c$ for all $s\in S$, and
  • $\rho(b)=\ell$, where $b$ is the $B$-value of $u$.
Lemma 3.4.
Let $S,C,B$ be finite sets with $|C|\geq 6$. If $\pi\in\mathrm{Alt}(C)$ and $c$ is in the support of $\pi$, then $\psi_{\pi,\ell,c}$ is a permutation, and $\psi_{\pi,\ell,c}\in\mathcal{G}_{S,C,B}$.
Proof.
Suppose first that $c$ is not in the support of $\pi$. Then it is easy to see that again the set of nodes where the $C$-value is changed is invariant under the action, and again the action of $\psi_{\pi,\ell,c}$ has finite order, thus is a permutation. Namely, $\psi_{\pi,\ell,c}$ does not change the nodes where the $C$-value is $c$ whether or not $\pi$ is applied (since $c$ is not in the support of $\pi$), and the other two conditions speak only about $B$-values, which are never affected.
Since $|C|\geq 6$, we can write any permutation $\pi$ without $c$ in its support as $\pi=[\pi_{1},\pi_{2},\pi_{3},\pi_{4}]$. We now show that
$\psi_{\pi,\ell,c}=[\phi_{\pi_{1},s,c},\gamma_{\pi_{2}},\beta_{\pi_{3},\ell}].$
For this, observe that none of the maps $\phi_{\pi_{1},s,c},\gamma_{\pi_{2}},\beta_{\pi_{3},\ell}$ affect the set of cells where the others act: The only dependence of any of them on $C$-values is that whether $\phi_{\pi_{1},s,c}$ applies the permutation $\pi_{1}$ depends on the set of nodes with $C$-value $c$. Since $c$ is not in the support of any permutation, so this set never changes.
This means that in any particular cell, some of the $\phi_{\pi_{1},s,c},\gamma_{\pi_{2}},\beta_{\pi_{3},\ell}$ apply whenever they appear in the commutator (with positive or negative power), and some never apply. If even one of them never applies, then the commutator cancels. If all of them always apply, then the permutation applied on the $C$-value is $[\pi_{1},\pi_{2},\pi_{3}]=\pi$.
This happens if and only if the conditions of all three maps hold, i.e. precisely if
  • each existing $us$ for $s\in S$ has $C$-value $c$ (the condition for $\phi_{\pi_{1},c}$);
  • $u$ is successful (the condition for $\gamma_{\pi_{2}}$);
  • $\rho(u)=\ell$ (the condition for $\beta_{\pi_{3},\ell}$).
This is precisely the definition of $\psi_{\pi,\ell,c}$.
Next, suppose $c$ is in the support of $\pi$. We again first explain why $\psi_{\pi,\ell,c}$ is at least a bijection. Since we only consider ranked graphs, if node $u$ is successful and has color $(c,b)$ with $\rho(b)=\ell$, then $us$ has $B$-component $bs$ in its color, with $\rho(bs)=\ell+1$. Therefore, the $C$-component of this node is not affected by $\psi_{\pi,\ell,c}$, and it is easy to see that $\psi_{\pi,\ell,c}$ is again of finite order, thus a permutation.
@@
[…]
Namely, if $\rho(u)\neq\ell$, then clearly neither automorphism acts on the $C$-component of the color at $u$.
[…]
Definition 3.5.
The cone $G(u)$ at the node $u$ in a graph $G\in\mathcal{Q}_{S,C,D}$ is the smallest set $N$ of such that $u\in N$, and if $v\in N$ is successful, then $vS\subset N$. The node $u$ is the root of the cone. Write $G_{i}(u)=G(u)\cap uS^{i}$. The cone of depth $n$ $G_{\leq n}(u)$ is $\bigcup_{j=0}^{n}G_{j}(u)$. We say the cone is full if $G_{\leq n}(u)=uS^{\leq n}$.
Note that the nodes in $G_{i}(u)$ need not themselves be successful.
Lemma 3.6 (Ripple catching lemma).
Consider the group $\mathcal{G}_{S,C,B}$ with $B,C$ finite sets with $|C|\geq 4$. Let $\pi\in\mathrm{Alt}(C)$. Then there is an element in $\mathcal{G}_{S,C,B}$ of word $n$-norm linear in $n$, whose behavior can be described as follows: in the $C$-component of a node $u$, we apply $\pi$ if and only if there is a full $n$-cone at $u$, and for all $u\neq v\in G_{\leq n}(u)$, the $C$-value is 00.
Proof.
[…]
$\psi_{1}=\left(\prod_{\ell=1}^{n-1}\psi_{\pi,\ell,1}\right)\circ\gamma_{\pi,n}$
and
$\psi_{2}=\left(\prod_{\ell=1}^{n-1}\psi_{\pi^{-1},\ell,2}\right)\circ\gamma_{\pi^{-1},n},$
(where we recall that the product puts the $i=0$ value first, so $\psi_{\pi,1,0}$ is applied last when we act from the left).
Now, let $\pi_{1},\pi_{2}$ be arbitrary permutations such that $[\pi_{1},\pi_{2}]\neq 1$, and consider
[…]
We observe first that $\psi_{\pi_{2},0,1}$ and $\psi_{\pi_{2},0,1}$ only affect the $C$-values of nodes $u$ where the $B$-value $b$ satisfies $\rho(b)=0$.
[…]
Consider now a successful node $u$ where the $B$-value $b$ satisfies $\rho(b)=0$. We claim that $[\pi_{2},\pi_{1}]$ is applied to the $C$-value if and only if all nodes $ut$ exist and are successful for $t\in S^{\ell}$ for $0\leq\ell<n$, and the $C$-value is 00 in all nodes $ut$ with $t\in S^{\ell}$ for $0<\ell<n$.
[…]
For $j\in\{1,\ldots,n-2\}$, we say $v\in G_{j}(u)$ is active in ripple $p\in\{1,2\}$ if after applying
$(\prod_{\ell=1}^{i-1}\psi_{\pi,\ell,p})^{-1}\psi_{p}$
(i.e. just after applying the first action that affects the nodes at the level of $v$), the $C$-value at $v$ is $p$. Note that $\psi$ acts nontrivially at $u$ if and only if all nodes in $uS$ are active in both ripples (note that $uS=G_{1}(u)$ by the assumption that $u$ is successful).
We now observe that if $v\in G_{j-1}(u)$ and $vs$ is not active in ripple $p$, then if $j>1$, the node $v$ can not be active in both of the ripples $p,3-p$. Namely, suppose $v$ is active in ripple $1$. Then necessarily we have either that
  • $v$ has $C$-value 00 and $vs$ is active in ripple $1$, or
  • $v$ has $C$-value $1$ and $vs$ is inactive in ripple $1$.
Similarly, if $v$ is active in ripple $2$, then either
  • $v$ has $C$-value 00 and $vs$ is active in ripple $2$, or
  • $v$ has $C$-value $2$ and $vs$ is inactive in ripple $2$.
If $v$ has $C$-value 00, then since $vs$ is not active in both ripples, $v$ will not be either. If the $C$-value of $v$ is $1$ or $2$ originally, then it has the same value when either ripple reaches it, so it cannot be $1$ during one ripple and $2$ in the other.
We conclude that the only way $\psi$ can modify the $C$-value at $u$ is indeed that all $v\in G_{j}(x)$ are active for $11\leq j\leq n-$ 1, which means precisely that $u$ starts a full valid cone full of 00s. ∎
As already suggested in the beginning of the section, when we apply this, we will typically encode graphs in configurations in various ways, and then look for cones that additionally satisfy some conditions (that are sufficiently quick to check). Often, the colors will actually be read in multiple nodes of another group (usually, the group $G$ on which configurations are laid).
In particular, in the case of complexity-theoretic results, by carefully choosing whether or not to include a node in the encoded graph, and choosing the meaning of “successfulness” so that it means that some type of computation is propagated, we can ensure that successful cones also satisfy some constraints, which will allow us to encode Turing machine computations on them.
4 Nonembedding results
We start with the nonembedding results, as they involve a simpler version of the construction.